# Compositional Deep Probabilistic Models of DNA Encoded Libraries

### v1

@@ Abstract
DNA-Encoded Library (DEL) has proven to be a powerful tool that utilizes combinatorially constructed small molecules to facilitate highly efficient screening assays. These selection experiments, involving multiple stages of washing, elution, and identification of potent binders via unique DNA barcodes, often generate complex data. This complexity can potentially mask the underlying signals, necessitating the application of computational tools such as machine learning to uncover valuable insights. We introduce a compositional deep probabilistic model of DEL data, DEL-Compose, which decomposes molecular representations into their mono-synthon, di-synthon, and tri-synthon building blocks and capitalizes on the inherent hierarchical structure of these molecules by modeling latent reactions between embedded synthons. Additionally, we investigate methods to improve the observation models for DEL count data such as integrating covariate factors to more effectively account for data noise. Across two popular public benchmark datasets (CA-IX and HRP), our model demonstrates strong performance compared to count baselines, enriches the correct pharmacophores, and offers valuable insights via its intrinsic interpretable structure, thereby providing a robust tool for the analysis of DEL data.
@@ 1 Introduction
Figure 1: DELs are synthesized by sequential addition of chemically compatible building blocks. The composition of each molecule and its order of synthesis is encoded in the DNA barcode. The pool of DEL molecules is then used in selection experiments that washes off any weak binders both in the presence of the protein of interest (target) and in the absence of the protein (matrix). The DNA of surviving library members are then amplified and sequenced to get count data correlated to likely binders.
@@
Many of these methods are predicated on computing an enrichment score for each molecule as a function of the observed data or molecule structure. Gerry et al. 9 computes this enrichment score by fitting Poisson distributions to the count data and then computing the ratio of the on-target versus off-target binding affinities derived from the fitted distributions. However, this approach and other similar approaches 10, 11 do not extend to out-of-library predictions as the enrichment is computed from the data itself rather than predicted via molecular structure. To that end, other methods have tackled this problem by utilizing molecule structure via molecular fingerprints and graph neural networks. McCloskey et al. 12 and Zhang et al. 13 bin the count data and construct a classification problem based on their discretizations of the data. In particular, Zhang et al. 13 proposes exploiting the compositional structure of DELs for the extraction of enrichment signals, but unlike our work, no explicit generative models of this factorized representation or the improved likelihoods to deal with DEL data are built. Other approaches formulate the problem as a latent-variable prediction task, maximizing the probability of observing the count data under some prescribed probability distribution such as the Poisson or Negative Binomial distribution 14, 15, 16. Shmilovich et al. 17 extends the representation capabilities of models on DEL data by incorporating 3-D docked poses to enhance the performance of models without requiring additional supervised validation data. However, these prior works do not leverage the inherent hierarchical structure of DEL data.
@@
which explicitly factorizes the molecular representation in a motivated manner through the construction of a generative model. We propose learning individual synthon representations, and construct the corresponding di-synthon and tri-synthon representation from their respective synthon compositions utilizing learned latent reactions parametrized by neural networks. Signals in DEL selection experiments are obfuscated by the various sources of noise, but given the combinatorial construction of the library, these sources of variation are highly correlated within any particular synthon group. Naturally, this decomposition allows us to better attribute interpretability in our model, and our method additionally avoids the necessity of enumerating full-molecule structures, which can often be a tedious process prone to errors.
@@
for two targets: Carbonic Anhyrase IX (CA-IX) and Horseradish peroxidase (HRP) 9. Since these two well-studied targets have known pharmacophores, we demonstrate that our model can effectively pick out the important synthons–even on challenging splits of the data. Furthermore, we demonstrate that our method can obtain competitive performance even without requiring fully enumerated molecule structures. Lastly, we show that our model has nice interpretability, offering useful insights into the predictions given by the model.
@@ 2 Methods
Figure 2: DELs are synthesized in a sequential manner via a split-and-pool method. The DNA barcode encode the identity of the synthon building block that should be added at each step.
@@ 2.1 Molecules As Synthon Compositions
Let $\mathcal{X}$ be the set of DEL molecules in our dataset, and $\{\mathcal{S}_{A},\mathcal{S}_{B},\mathcal{S}_{C}\}$ be the sets of synthons at the first, second and third positions respectively. Each molecule is denoted by $x_{abc}\in\mathcal{X}$, where the subscript indicates the identity of the synthon at a particular position ($a\in\mathcal{S}_{A},b\in\mathcal{S}_{B},c\in\mathcal{S}_{C}$). To simplify (and overload) notation, we omit the subscript for a particular synthon position if it is absent. For instance, $x_{b}$ denotes the molecule corresponding to the synthon $b$ at the second position, and $x_{ab}$ denotes the molecule corresponding to the combination of synthon $a$ at the first position and $b$ at the second position. We can generalize this process trivially to higher order synthon composition over more than 3 synthons, but for the purpose of presentation, we focus on the setting of tri-synthon DELs in this work.
@@ 2.2 DEL-Compose: A Generative Model of DEL Data
Here we introduce DEL-Compose, a general paradigm for modeling DEL molecules leveraging their combinatorial nature (Figure 3). Our objective is to maximize the likelihood of observing the count data given an input molecule $x_{abc}$. Let $\mathcal{Z}=\{z_{s}\in\mathcal{R}^{d}|s\in[a,b,c,ab,abc,mol]\}$ be a collection of latent synthon embeddings each of dimension $d$. $z_{a}$ denotes an embedding of an individual synthon $x_{a}$, while $z_{ab}$ denotes an embedding of a di-synthon $x_{ab}$ which is the reaction output of $x_{a}$ and $x_{b}$. Similarly, $z_{abc}$ denotes an embedding of tri-synthon $x_{abc}$, which itself is a product of the reaction of $x_{ab}$ and $x_{c}$. Finally, $z_{mol}$ is the aggregated embedding of all the above representations. We thus assume embeddings corresponding to all the (partial) products and building blocks of molecule can be viewed as a synthon decomposition and their respective set of chemical reactions.
@@ 3 Data
Figure 5: Known pharmacophores for both CA-IX (top) and HRP (bottom) pictured in descending order of affinity. For CA-IX, benzene-sulfonamides are known structures to induce affinity. The substitution of the sulfonamide affects the reactivity of the chemical specie, wherein the para-substituted constituent is found to be much more active. For HRP, electrophilic Michael acceptors are known pharmacophores.
@@ 4 Results and Discussion
For all results we present in the following, models are trained to minimize the negative log likelihood (NLL): $-\log[p(C_{t},C_{m},\Phi_{mol},\mathcal{Z}|x_{a},x_{b},x_{c};\Theta)]$ through gradient descent on parameters $\Theta$ using the Adam optimizer.
@@ 4.2 Setup for out-of-distribution generalization
Random splits are not always ideal for testing molecule datasets 22, 23. In order to test the generalizability of molecule representations, many approaches attempt to split molecule by molecular scaffolds 24. For DELs, rather than generic molecular scaffolding strategies, synthons provide a natural grouping and separation of the chemical space. By using synthons to split the data, we can test the generalizability of the model on unseen chemical structures. In the dataset that we are using, the known pharmacophores are conveniently localized to specific synthons, so we can develop intuitive splitting strategies. Most of the signal is captured by these pharmacophores, so we cannot withhold all of these molecules from training. Instead, we split on the synthon position that does not include these individual pharmacophores. Specifically, for CA-IX, the benzene-sulfonamides are at the C position, so we create synthon splits by splitting on the B position. For HRP, the electrophilic Michael acceptors are at the B position, so we split the data at the C position.
@@ 4.5 DEL-Compose captures enrichment of important pharmacophores
Next, we compare our deep probabilistic approach to several baselines that compute enrichments based on counts alone. These enrichment baselines try to assess the affinity of a molecule based on some assumed functional form. For instance, the Diff Enrichment score makes the assumption that there is a simple additive effect between the matrix and target counts. Poisson enrichment is taken from Gerry et al. 9, which computes a maximum likelihood Poisson distribution for the target and matrix counts and then computes a ratio of the target at the lower 95% confidence interval (CI) and the matrix at the upper 95% CI. This baseline is perhaps closest to our model, however, our model relates the two distributions directly, whereas this score does not directly correlate the two values.
@@
Random Split Matrix NLL $\downarrow$ Target NLL $\downarrow$ Sum NLL $\downarrow$ PRC-AUC $\uparrow$ CA-IX Diff Enrichment - - - 0.23 $\pm$ 0.01 Ratio Enrichment - - - 0.26 $\pm$ 0.02 Poisson Enrichment - - - 0.25 $\pm$ 0.01 DEL-Compose 3.17 $\pm$ 0.03 2.82 $\pm$ 0.04 5.99 $\pm$ 0.06 0.28 $\pm$ 0.33 DEL-Compose (Pre) 2.97 $\pm$ 0.01 2.80 $\pm$ 0.02 5.77 $\pm$ 0.02 0.90 $\pm$ 0.03 DEL-Compose (Rep) 3.13 $\pm$ 0.03 2.65 $\pm$ 0.03 5.78 $\pm$ 0.06 0.73 $\pm$ 0.34 DEL-Compose (Pre+Rep) 2.96 $\pm$ 0.02 2.65 $\pm$ 0.01 5.61 $\pm$ 0.01 0.84 $\pm$ 0.07 HRP Diff Enrichment - - - 0.59 $\pm$ 0.01 Ratio Enrichment - - - 0.48 $\pm$ 0.01 Poisson Enrichment - - - 0.57 $\pm$ 0.01 DEL-Compose 6.51 $\pm$ 0.11 5.61 $\pm$ 0.09 12.12 $\pm$ 0.19 0.80 $\pm$ 0.12 DEL-Compose (Pre) 6.30 $\pm$ 0.03 5.35 $\pm$ 0.02 11.65 $\pm$ 0.04 0.78 $\pm$ 0.03 DEL-Compose (Rep) 6.39 $\pm$ 0.04 5.53 $\pm$ 0.04 11.92 $\pm$ 0.08 0.81 $\pm$ 0.04 DEL-Compose (Pre+Rep) 6.23 $\pm$ 0.02 5.30 $\pm$ 0.02 11.54 $\pm$ 0.03 0.80 $\pm$ 0.04
Table 1: Metrics for different variants of DEL-Compose compared to baseline enrichment metrics on random splits. Metrics are averaged across the test set over 5 different splits.
Synthon Split Matrix NLL $\downarrow$ Target NLL $\downarrow$ Sum NLL $\downarrow$ PRC-AUC $\uparrow$ CA-IX Diff Enrichment - - - 0.22 $\pm$ 0.02 Ratio Enrichment - - - 0.25 $\pm$ 0.02 Poisson Enrichment - - - 0.24 $\pm$ 0.02 DEL-Compose 3.58 $\pm$ 0.34 2.83 $\pm$ 0.17 6.41 $\pm$ 0.46 0.13 $\pm$ 0.01 DEL-Compose (Pre) 3.13 $\pm$ 0.13 2.81 $\pm$ 0.18 5.94 $\pm$ 0.31 0.63 $\pm$ 0.42 DEL-Compose (Rep) 3.54 $\pm$ 0.29 2.64 $\pm$ 0.14 6.19 $\pm$ 0.39 0.73 $\pm$ 0.33 DEL-Compose (Pre+Rep) 3.11 $\pm$ 0.11 2.61 $\pm$ 0.14 5.72 $\pm$ 0.25 0.75 $\pm$ 0.33 HRP Diff Enrichment - - - 0.59 $\pm$ 0.01 Ratio Enrichment - - - 0.48 $\pm$ 0.01 Poisson Enrichment - - - 0.56 $\pm$ 0.01 DEL-Compose 8.09 $\pm$ 2.04 7.36 $\pm$ 1.28 15.45 $\pm$ 3.32 0.85 $\pm$ 0.16 DEL-Compose (Pre) 7.44 $\pm$ 2.09 6.46 $\pm$ 1.35 13.90 $\pm$ 3.44 0.85 $\pm$ 0.13 DEL-Compose (Rep) 8.75 $\pm$ 3.73 7.36 $\pm$ 1.69 16.11 $\pm$ 5.41 0.88 $\pm$ 0.05 DEL-Compose (Pre+Rep) 7.29 $\pm$ 2.11 6.26 $\pm$ 1.26 13.55 $\pm$ 3.36 0.90 $\pm$ 0.06
Table 2: Metrics for different variants of DEL-Compose compared to baseline enrichment metrics on synthon splits. Compared to random splits, the average NLL loss is higher, which confirms our belief that these are more challenging splits of the data for a model to learn over. Our model still outperforms baselines and ablations even on these more challenging splits.

### v2

# Compositional Deep Probabilistic Models of DNA-Encoded Libraries

@@ Abstract
DNA-Encoded Library (DEL) has proven to be a powerful tool that utilizes combinatorially constructed small molecules to facilitate highly efficient screening experiments. These selection experiments, involving multiple stages of washing, elution, and identification of potent binders via unique DNA barcodes, often generate complex data. This complexity can potentially mask the underlying signals, necessitating the application of computational tools such as machine learning to uncover valuable insights. We introduce a compositional deep probabilistic model of DEL data, DEL-Compose, which decomposes molecular representations into their mono-synthon, di-synthon, and tri-synthon building blocks and capitalizes on the inherent hierarchical structure of these molecules by modeling latent reactions between embedded synthons. Additionally, we investigate methods to improve the observation models for DEL count data such as integrating covariate factors to more effectively account for data noise. Across two popular public benchmark datasets (CA-IX and HRP), our model demonstrates strong performance compared to count baselines, enriches the correct pharmacophores, and offers valuable insights via its intrinsic interpretable structure, thereby providing a robust tool for the analysis of DEL data.
@@ 1 Introduction
Many of these methods are predicated on computing an enrichment score for each molecule as a function of the observed data or molecule structure. Gerry et al. 9 computes this enrichment score by fitting Poisson distributions to the count data and then computing the ratio of the on-target versus off-target binding events derived from the fitted distributions. However, this approach and other similar approaches 10, 11 do not extend to out-of-library predictions as the enrichment is computed from the data itself rather than predicted via molecular structure. To that end, other methods have tackled this problem by utilizing molecule structure via molecular fingerprints and graph neural networks. McCloskey et al. 12 and Zhang et al. 13 bin the count data and construct a classification problem based on their discretizations of the data. In particular, Zhang et al. 13 proposes exploiting the compositional structure of DELs for the extraction of enrichment signals, but unlike our work, no explicit generative models of this factorized representation or the improved likelihoods to deal with DEL data are built. Other approaches formulate the problem as a latent-variable prediction task, maximizing the probability of observing the count data under some prescribed probability distribution such as the Poisson or Negative Binomial distribution 14, 15, 16. Shmilovich et al. 17 extends the representation capabilities of models on DEL data by incorporating 3-D docked poses to enhance the performance of models without requiring additional supervised validation data. However, these prior works do not leverage the inherent hierarchical structure of DEL data.
Figure 1: a) DELs are synthesized in a sequential manner via a split-and-pool method. The DNA barcode encode the identity of the synthon building block that should be added at each step. b) The pool of DEL molecules is then used in selection experiments that washes off any weak binders both in the presence of the protein of interest (target) and in the absence of the protein (matrix). The DNA of surviving library members are then amplified and sequenced to get count data correlated to likely binders.
@@
which explicitly factorizes the molecular representation in a motivated manner through the construction of a generative model. We propose learning individual synthon representations, and construct the corresponding di-synthon and tri-synthon representation from their respective synthon composition parametrized by neural networks. We use the term latent reactions to denote the construction of these embeddings, which is similar to how molecules are synthesized via chemical reactions, but in embedding space. Signals in DEL selection experiments are obfuscated by the various sources of noise, but given the combinatorial construction of the library, these sources of variation are highly correlated within any particular synthon group. For instance, PCR bias which can arise from specific codons will result in correlated biases for a particular synthon. Naturally, this decomposition allows us to better attribute interpretability in our model, and our method additionally avoids the necessity of enumerating full-molecule structures, which requires careful specification of reaction templates which can be tedious and result in errors.
@@
for two targets: Carbonic Anhyrase IX (CA-IX) and Horseradish peroxidase (HRP) 9. Since these two well-studied targets have known pharmacophores, we demonstrate that our model can effectively pick out the important synthons–even on challenging splits of the data. Furthermore, we demonstrate that our method can obtain competitive performance even without requiring fully enumerated molecule structures. Lastly, we show that our model offers useful insights into the predictions given by the model by assigning importance weights to synthons combinations via attention mechanisms, which correctly ranks synthon importance, and predicting possible dropout in the data to illustrate inferred noise.
@@ 2.1 Molecules As Synthon Compositions
Let $\mathcal{X}$ be the set of DEL molecules in our dataset, and $\{\mathcal{S}_{A},\mathcal{S}_{B},\mathcal{S}_{C}\}$ be the sets of synthons at the first, second and third positions respectively. Each molecule is denoted by $x_{abc}\in\mathcal{X}$, where the subscript indicates the identity of the synthon at a particular position ($a\in\mathcal{S}_{A},b\in\mathcal{S}_{B},c\in\mathcal{S}_{C}$). To simplify (and overload) notation, we omit the subscript for a particular synthon position if it is absent. For instance, $x_{b}$ denotes the molecule corresponding to the synthon $b$ at the second position, and $x_{ab}$ denotes the molecule corresponding to the combination of synthon $a$ at the first position and $b$ at the second position.
@@ 2.2 DEL-Compose: A Generative Model of DEL Data
Here we introduce DEL-Compose, a general paradigm for modeling DEL molecules leveraging their combinatorial nature (Figure 2), and we formulate the objective for a typical DEL design, but recognize that other variants exist. Our objective is to maximize the likelihood of observing the count data given an input molecule $x_{abc}$. Let $\mathcal{Z}=\{z_{s}\in\mathcal{R}^{d}|s\in[a,b,c,ab,abc,mol]\}$ be a collection of latent synthon embeddings each of dimension $d$. $z_{a}$ denotes an embedding of an individual synthon $x_{a}$, while $z_{ab}$ denotes an embedding of a di-synthon $x_{ab}$ which is the reaction output of $x_{a}$ and $x_{b}$. Similarly, $z_{abc}$ denotes an embedding of tri-synthon $x_{abc}$, which itself is a product of the reaction of $x_{ab}$ and $x_{c}$. Finally, $z_{mol}$ is the aggregated embedding of all the above representations. We thus assume embeddings corresponding to all the (partial) products and building blocks of molecule can be viewed as a synthon decomposition and their respective set of chemical reactions.
@@ 3 Data
Figure 4: Known pharmacophores for both CA-IX (top) and HRP (bottom) pictured in descending order of affinity. For CA-IX, benzene-sulfonamides are known structures to induce affinity. The substitution of the sulfonamide affects the reactivity of the chemical specie, wherein the para-substituted constituent is found to be much more active. For HRP, electrophilic Michael acceptors are known pharmacophores. The structures are labeled with attachment to other synthons: for CA-IX, the sulfonamides are at the C positon with attachment to synthon B, and for HRP, the Michael acceptors are at the B position, with attachment to synthon A.
@@ 4 Results and Discussion
For all results we present in the following, models are trained to minimize the negative log likelihood (NLL): $-\log[p(C_{t},C_{m},\Phi_{mol},\mathcal{Z}|x_{a},x_{b},x_{c};\Theta)]$ through gradient descent on parameters $\Theta$ using the Adam optimizer  22, which is a typical optimization technique that utilizes moving averages of gradients to help the model converge faster.
@@ 4.2 Setup for out-of-distribution generalization
Random splits are not always ideal for testing molecule datasets 23, 24. In order to test the generalizability of molecule representations, many approaches attempt to split molecule by molecular scaffolds 25. For DELs, rather than generic molecular scaffolding strategies, synthons provide a natural grouping and separation of the chemical space. By using synthons to split the data, we can test the generalizability of the model on unseen chemical structures. In the dataset that we are using, the known pharmacophores are conveniently localized to specific synthons, so we can develop intuitive splitting strategies. Most of the signal is captured by these pharmacophores, so we cannot withhold all of these molecules from training. Instead, we split on the synthon position that does not include these individual pharmacophores. Specifically, for CA-IX, the benzene-sulfonamides are at the C position, so we create synthon splits by splitting on the B position. For HRP, the electrophilic Michael acceptors are at the B position, so we split the data at the C position.
@@ 4.5 DEL-Compose captures enrichment of important pharmacophores
Next, we compare our deep probabilistic approach to several baselines that compute enrichments based on counts alone. These enrichment baselines try to assess the affinity of a molecule based on some assumed functional form. For instance, the Diff Enrichment score makes the assumption that there is a simple additive effect between the matrix and target counts. Poisson enrichment is taken from Gerry et al. 9, which computes a maximum likelihood Poisson distribution for the target and matrix counts and then computes a ratio of the target at the lower 95% confidence interval (CI) and the matrix at the upper 95% CI. This baseline is perhaps closest to our model, however, our model relates the two distributions directly, whereas this score does not directly correlate the two values. Additionally, we also include the deldenoiser model 26, which computes a fitness value for each count that correlates to the denoised signal of the counts. We run their model using default parameters, without yield data since there is no yield data for these datasets, and compute enrichment as the average fitness across each target replicate.
@@
Random Split Matrix NLL $\downarrow$ Target NLL $\downarrow$ Sum NLL $\downarrow$ PRC-AUC $\uparrow$ CA-IX Diff Enrichment - - - 0.23 $\pm$ 0.01 Ratio Enrichment - - - 0.26 $\pm$ 0.02 Poisson Enrichment - - - 0.25 $\pm$ 0.01 deldenoiser - - - 0.19 $\pm$ 0.01 DEL-Compose 3.17 $\pm$ 0.03 2.82 $\pm$ 0.04 5.99 $\pm$ 0.06 0.28 $\pm$ 0.33 DEL-Compose (Pre) 2.97 $\pm$ 0.01 2.80 $\pm$ 0.02 5.77 $\pm$ 0.02 0.90 $\pm$ 0.03 DEL-Compose (Rep) 3.13 $\pm$ 0.03 2.65 $\pm$ 0.03 5.78 $\pm$ 0.06 0.73 $\pm$ 0.34 DEL-Compose (Pre+Rep) 2.96 $\pm$ 0.02 2.65 $\pm$ 0.01 5.61 $\pm$ 0.01 0.84 $\pm$ 0.07 HRP Diff Enrichment - - - 0.59 $\pm$ 0.01 Ratio Enrichment - - - 0.48 $\pm$ 0.01 Poisson Enrichment - - - 0.57 $\pm$ 0.01 deldenoiser - - - 0.26 $\pm$ 0.01 DEL-Compose 6.51 $\pm$ 0.11 5.61 $\pm$ 0.09 12.12 $\pm$ 0.19 0.80 $\pm$ 0.12 DEL-Compose (Pre) 6.30 $\pm$ 0.03 5.35 $\pm$ 0.02 11.65 $\pm$ 0.04 0.78 $\pm$ 0.03 DEL-Compose (Rep) 6.39 $\pm$ 0.04 5.53 $\pm$ 0.04 11.92 $\pm$ 0.08 0.81 $\pm$ 0.04 DEL-Compose (Pre+Rep) 6.23 $\pm$ 0.02 5.30 $\pm$ 0.02 11.54 $\pm$ 0.03 0.80 $\pm$ 0.04
Table 1: Metrics for different variants of DEL-Compose compared to baselines on random splits. Metrics are averaged across the test set over 5 different splits.
Synthon Split Matrix NLL $\downarrow$ Target NLL $\downarrow$ Sum NLL $\downarrow$ PRC-AUC $\uparrow$ CA-IX Diff Enrichment - - - 0.22 $\pm$ 0.02 Ratio Enrichment - - - 0.25 $\pm$ 0.02 Poisson Enrichment - - - 0.24 $\pm$ 0.02 deldenoiser - - - 0.19 $\pm$ 0.00 DEL-Compose 3.58 $\pm$ 0.34 2.83 $\pm$ 0.17 6.41 $\pm$ 0.46 0.13 $\pm$ 0.01 DEL-Compose (Pre) 3.13 $\pm$ 0.13 2.81 $\pm$ 0.18 5.94 $\pm$ 0.31 0.63 $\pm$ 0.42 DEL-Compose (Rep) 3.54 $\pm$ 0.29 2.64 $\pm$ 0.14 6.19 $\pm$ 0.39 0.73 $\pm$ 0.33 DEL-Compose (Pre+Rep) 3.11 $\pm$ 0.11 2.61 $\pm$ 0.14 5.72 $\pm$ 0.25 0.75 $\pm$ 0.33 HRP Diff Enrichment - - - 0.59 $\pm$ 0.01 Ratio Enrichment - - - 0.48 $\pm$ 0.01 Poisson Enrichment - - - 0.56 $\pm$ 0.01 deldenoiser - - - 0.27 $\pm$ 0.00 DEL-Compose 8.09 $\pm$ 2.04 7.36 $\pm$ 1.28 15.45 $\pm$ 3.32 0.85 $\pm$ 0.16 DEL-Compose (Pre) 7.44 $\pm$ 2.09 6.46 $\pm$ 1.35 13.90 $\pm$ 3.44 0.85 $\pm$ 0.13 DEL-Compose (Rep) 8.75 $\pm$ 3.73 7.36 $\pm$ 1.69 16.11 $\pm$ 5.41 0.88 $\pm$ 0.05 DEL-Compose (Pre+Rep) 7.29 $\pm$ 2.11 6.26 $\pm$ 1.26 13.55 $\pm$ 3.36 0.90 $\pm$ 0.06
Table 2: Metrics for different variants of DEL-Compose compared to baselines on synthon splits. Compared to random splits, the average NLL loss is higher, which confirms our belief that these are more challenging splits of the data for a model to learn over. Our model still outperforms baselines and ablations even on these more challenging splits.